# DNAMotifTokenizer: Towards Biologically Informed Tokenization of Genomic Sequences

### v1

@@ 1 Introduction
Recent advances in artificial intelligence (AI) and large language models (LLMs) have transformed nearly every field of biological research. By analyzing complex, noisy, and large-scale datasets, these models can uncover hidden patterns, generate predictions, and accelerate the discovery of new biological knowledge and molecular structures (embedding_ai_biology_2024). In genetics, building on the remarkable success of text-based LLMs, researchers have developed DNA language models (DNA-LMs) to capture the latent “grammar” of genomic sequences. These models are being leveraged to improve DNA sequence design, investigate the genetic basis of evolution, and interpret genetic mutations underlying human traits and diseases.
Over the past few years, a series of DNA-LMs has emerged (see Figure 1a). Early efforts, such as DNABERT-1 (ji2021dnabert), introduced k-mer–based tokenizers and transformer architectures to model DNA sequences, laying the foundation for various downstream applications. DNABERT-2 (zhou2023dnabert) extended this idea by introducing byte-pair encoding (BPE)-based (sennrich2015neural) tokenizer and pretrained on multi-species genomes. A large-scale model, Nucleotide Transformer (dalla2025nucleotide), has scaled up in both parameter size and training corpus, improving accuracy and generalizability. The HyenaDNA (nguyen2023hyenadna) has explored long-context modeling to better capture distal dependencies in the genome. More recently, Evo-2 (brixi2025genome) has been developed to expand prediction and design across DNA, RNA, and proteins. Collectively, these models underscore both the promise and challenges of scaling DNA-LMs for biological discovery, including regulatory element prediction, non-coding genetic variant interpretation, and DNA sequence designs.
Despite their superb fine-tuning performance in downstream tasks, current DNA-LMs often exhibit poor zero-shot generalization to new tasks (patel2024dart). The bottleneck lies largely in the DNA tokenization process, which breaks down raw DNA sequences into fundamental units for the model to process (see Figure 1b). Standard tokenization strategies, such as fixed k-mers or subword methods like byte-pair encoding (BPE), often fail to efficiently capture these biologically meaningful DNA motifs. It is critical to optimize the DNA tokenization step towards the development of accurate, interpretable, and generalizable models.
@@ 2 Background
The k-mer and BPE-based tokenizers are commonly used in state-of-the-art DNA-LMs(Table 1). HyenaDNA (nguyen2023hyenadna) uses a 1-mer tokenizer and a decoder-only architecture with the Hyena operator to model long-range dependencies. DNABERT-1 (ji2021dnabert) tokenized the genome with overlapping k-mers and trained separate models for each, substantially outperforming several downstream tasks. Nucleotide Transformer (dalla2025nucleotide) employs non-overlapping 6-mer tokenization and leverages large-scale pretraining across thousands of human individual genomes, and hundreds of species. zhou2023dnabert proposed DNABERT-2, which uses a byte-pair encoding (BPE) tokenizer inspired by NLP. MxDNA (qiao2024model) introduced a learnable tokenizer using a Mixture-of-Experts framework (shazeer2017outrageously) and deformable convolutions (dai2017deformable).
@@
In addition, DNA is composed of highly repetitive, short, sparse, unevenly distributed, but conserved DNA sequence motifs across 600 million years of bilaterian evolution nitta2015conservation, largely represented by transcription factor (TF) binding motifs. The complexity of gene regulation often arises from the specific rearrangement and combination of these conserved motifs into context-specific regulatory elements wong2020deep. Standard tokenization methods, which are agnostic to biological function, often arbitrarily split these meaningful motifs into smaller, non-functional tokens (see Figure 1b). As a result, it hampers DNA-LMs’ ability to learn biologically meaningful representations of genomic sequences and complicates downstream model interpretation.
@@ 3.1 Genomic sequences
Human reference genome: We use the most widely used human reference genome, version hg38 (GRCh38) (grch38), which incorporates improved accuracy and coverage over previous releases.
Annotation of motif regions: We download the genome-wide JASPAR CORE TF motif predictions on the human reference genome (hg38) from the UCSC Genome Browser (lee2020ucsc)(raney2014track). For BPE training, we extract all predicted motif sequences and merge any overlapping ones to create a non-redundant set. The resulting set of motif sequences covers approximately 59.84% (See Table C.1.1) of the human genome.
Annotation of cCREs: Candidate cis-regulatory elements (cCREs) are functional regulatory units in the genome, such as promoters and enhancers, that are typically hundreds of base pairs long. These regions often contain multiple TF motifs and are crucial for controlling when and where genes are expressed. For this study, we download human cCRE annotations from The Encyclopedia of DNA Elements(ENCODE), which provide genomic coordinates and regulatory classifications (encode2020expanded)(luo2020new). We extract the corresponding DNA sequences for BPE training, which constitute approximately 20.32% (See Table C.2.2) of the human reference genome.
@@ 3.2 Benchmark datasets
Genome Understanding Evaluation(GUE): Genome Understanding Evaluation (GUE) dataset was collected by (zhou2023dnabert), consisting of 28 distinct datasets across 7 tasks and 4 species, with DNA inputs ranging from 70 to 1000 base pairs (bp). The metric we use is the Matthews Correlation Coefficient (MCC) (see Appendix).
Nucleotide Transformer Benchmarks: Nucleotide Transformer (NT) benchmarks were collected by (dalla2025nucleotide), it includes 18 datasets across 4 tasks only on humans. For this dataset, MCC is employed as the metric.
Dart-Eval: This dataset was introduced by (patel2024dart) and contains five tasks. In our experiments, we focus on tasks 1–3. Accuracy (ACC) served as the evaluation metric. Task 1 involves distinguishing cCRE regions from background sequences. Task 2 requires identifying transcription factor (TF) binding motifs within background sequences. Task 3 entails classifying sequences specific to five different cell types against background sequences.
Genomic Benchmarks: This dataset was collected by (grevsova2023genomic), it includes 9 tasks, across 9 species. We only use human related datasets. MCC is used as the metric.
SCREEN: We create this benchmark dataset by first downloading the cCREs on hg38 from the SCREEN interface (encode2020expanded), a platform for searching and visualizing the ENCODE Registry of candidate cis-Regulatory Elements (cCREs). This Registry contains 2,348,854 human cCREs, classified into eight categories, including promoter-like signatures (PLS), proximal and distal enhancer-like signatures (pELS, dELS), and CTCF- or TF–associated accessible elements (CA, CA-CTCF, CA-H3K4me3, CA-TF, TF). We generated a negative superset by taking the complement of all cCRE regions and dividing it into 300 base pair (bp) segments. For each cCRE category, we then randomly sample from this superset to obtain the same number of sequences as in the corresponding negative set. This procedure yielded eight datasets, each containing an equal number of positive and negative sequences.
@@ 4.1 Tokenizers
K-mer tokenizer: A k-mer is a substring of length k from a DNA or RNA sequence (moeckel2024survey). K-mers can be generated in two ways: overlapping, by sliding a window one base at a time to capture all subsequences, or non-overlapping, by moving the window in steps of k to produce disjoint subsequences. In our experiments, we test overlapping 3-mer and 6-mer implemented by DNABERT-1 (ji2021dnabert), and non-overlapping 6-mer implemented by Nucleotide Transformer (dalla2025nucleotide), respectively.
BPE tokenizer: BPE is a learnable subword tokenization algorithm that iteratively merges the most frequent pairs of characters or subwords in a training corpus to build a vocabulary of common subwords (sennrich2015neural). The resulting merge rules, based on pair frequency, are recorded and applied to tokenize new sequences consistently, allowing the model to capture recurring patterns efficiently. The BPE tokenizer for DNA sequences was introduced by DNABERT-2. We use the BPE tokenizer pretrained by DNABERT-2 and also train our own BPEs on human reference genome, motif enriched genomic regions, and cis-regulatory element (cCRE) regions (see Section 3). Three vocabulary sizes were explored: 4096 (matching DNABERT-2), 2048, and 1024. Each BPE was initialized with the DNA alphabet (A, T, C, G, N) and five special tokens ([PAD], [UNK], [CLS], [SEP], [MASK]), with a minimum frequency threshold of 100.
Others: MxDNA (qiao2024model) is excluded from comparison due to a lack of publicly available pre-trained weights.
@@ 4.2 Models
Architecture: We adopt a BERT-based masked language model architecture (BertMLM) (devlin2019bert) for our experiments. The model uses the Transformer encoder architecture (vaswani2017attention) with 12 layers and 12 attention heads and a hidden size of 768, resulting in intermediate feed-forward layers of size 3072. The maximum input sequence length is 512 tokens, and the model uses learnable positional embeddings up to this length. Dropout srivastava2014dropout is applied to both the attention probabilities and hidden layers with a rate of 0.1, and the GELU (hendrycks2016gaussian) activation function is used throughout. The model vocabulary size is changed according to the tokenizer used, and a type vocabulary of size 2 is included to distinguish segment embeddings. All parameters are initialized with a standard deviation of 0.02.
@@
We compare the vocabulary similarities by calculating the pairwise Jaccard Index. We also generate Venn diagrams illustrating their word overlap (hulsen2021biovenn). See Appendix for more details.
@@ 5 Benchmarking of State-of-the-art Tokenizers
We use the NT-benchmarks to investigate the zero-shot ability, where k-mer tokenizers consistently outperformed BPE tokenizers. Both overlapping and non-overlapping k-mer tokenizers struggled to cluster four distinct genomic elements, while BPE tokenizers performed better at separating enhancers and promoters from other DNA sequences (Figure 2c). Furthermore, we leverage the most comprehensive GUE benchmark and group downstream tasks by the length of DNA sequences. We find that k-mer tokenizers performed marginally better on shorter sequences (100 bp), whereas BPE tokenizers excel on longer ones (>>500 bp) (Figure 2d), which is consistent with previous work (zhou2023dnabert). Taken together, we conclude that BPE-based tokenizers achieve more robust performance on both zero-shot and fine-tuning tasks and are more generalizable for predictions involving longer DNA sequences.
@@ 7.1 Algorithm
Motif processing: Motifs are short, recurring patterns in DNA, RNA, or protein sequences that are frequently associated with specific biological functions. In this work, we focus on TF motifs, which are short DNA sequences that are bound by transcription factor proteins to regulate gene expression. The TF motifs are generally represented in the form of position weight matrices (PWM)(stormo2000dna), which indicate the probability of each nucleotide occurring at each position within the motif. We download motif PWMs from the JASPAR 2024 motif library (sandelin2004jaspar)(rauluseviciute2024jaspar), which is a widely used, open-access repository. We use the vertebrate library, which contains 879 non-redundant motifs. As most TF motifs range in length from 5 to 12 base pairs (bp), we exclude any motifs longer than 12 bp from our analysis. To incorporate TF motifs into our vocabulary, we binarize their PWMs and encode them into fixed sequences (Figure  5a). For each TF PWM, we apply a threshold probability of 0.5. Positions with lower probability are defined as wildcard positions. Subsequently, we discard wildcard positions at both ends of the motif. For the remaining positions, we encode them using the nucleotide with the highest probability.
@@ 7.2 Evaluation performance
Figure 6: Interpretation of DNAMotifTokenizer. (a) Workflow for the multiclass cell-type ATAC-seq classification task. (b) Distribution of average attribution scores for k-mer and motif tokens on the test set (10%). The motif tokens exhibit the strongest influence on model predictions. (c) Overlap between top 200 highest-attribution tokens from three cell types. (d) Highly attributed motif tokens matched with motif enrichment results from  li2023comparative.
Interpretability: We collected single-nucleus ATAC-seq (snATAC-seq) data generated from three diverse brain cell types: intratelencephalic neurons from cortical layer 2/3 (ITL23), VIP-positive GABAergic neurons (VIP), and Microglia (MGC) li2023comparative. We identified cell-type-specific ATAC-seq peak regions, which serve as prediction targets (Figure  6a). To control the vocabulary size and achieve fair comparison, we then fine-tuned our DNAMotifTokenizer models on this dataset. Then, we performed token attribution analysis using Integrated Gradients sundararajan2017axiomatic on the test set (10%). For each token, we computed its attribution score toward the true class label and then averaged the scores; Tokens with attribution scores farther away from zero—either positive or negative—indicate stronger influence on the model’s prediction (Figure  6b). By ranking them, we found that in our DNAMotifTokenizer model, most of the highly influential tokens correspond to motif tokens (in red), whereas the k-mer tokens (in blue) generally had attribution scores close to zero. We selected the top 200 most contributive motif tokens for each cell type. The Venn plot indicates that DNAMotifTokenizer not only utilizes 143 motif tokens shared between three cell types, but also uses 25, 26, 34 cell-type-specific motif tokens for prediction in ITL23, VIP, MGC, respectively (Figure  6c). In addition, we compared these most contributive motif tokens with the enriched motifs for each cell type reported in the original paper li2023comparative. We showed top-matched motifs from three cell types as a heatmap (Figure  6d). In ITL23 cell type, motifs from the WT1, RORa, Egr, and KLF families are captured and enriched, whereas only WT1, and Egr2 are captured in the VIP cell type. In MGC cell type, SOX family motifs are captured and enriched. These results demonstrate the interpretability of motif tokens introduced by DNAMotifTokenizer in various cell types.

### v2

@@ 1 Introduction
Recent advances in artificial intelligence (AI) and large language models (LLMs) have transformed nearly every field of biological research. By analyzing complex, noisy, and large-scale datasets, these models can uncover hidden patterns, generate predictions, and accelerate the discovery of new biological knowledge and molecular structures (Nature Methods, 2024). In genetics, building on the remarkable success of text-based LLMs, researchers have developed DNA language models (DNA-LMs) to capture the latent “grammar” of genomic sequences. These models are being leveraged to improve DNA sequence design, investigate the genetic basis of evolution, and interpret genetic mutations underlying human traits and diseases.
Over the past few years, a series of DNA-LMs has emerged (see Figure 1a). Early efforts, such as DNABERT-1 (Ji et al., 2021), introduced k-mer–based tokenizers and transformer architectures to model DNA sequences, laying the foundation for various downstream applications. DNABERT-2 (Zhou et al., 2023) extended this idea by introducing byte-pair encoding (BPE)-based (Sennrich et al., 2015) tokenizer and pretrained on multi-species genomes. A large-scale model, Nucleotide Transformer (Dalla-Torre et al., 2025), has scaled up in both parameter size and training corpus, improving accuracy and generalizability. The HyenaDNA (Nguyen et al., 2023) has explored long-context modeling to better capture distal dependencies in the genome. More recently, Evo-2 (Brixi et al., 2025) has been developed to expand prediction and design across DNA, RNA, and proteins. Collectively, these models underscore both the promise and challenges of scaling DNA-LMs for biological discovery, including regulatory element prediction, non-coding genetic variant interpretation, and DNA sequence designs.
Despite their superb fine-tuning performance in downstream tasks, current DNA-LMs often exhibit poor zero-shot generalization to new tasks (Patel et al., 2024). The bottleneck lies largely in the DNA tokenization process, which breaks down raw DNA sequences into fundamental units for the model to process (see Figure 1b). Standard tokenization strategies, such as fixed k-mers or subword methods like byte-pair encoding (BPE), often fail to efficiently capture these biologically meaningful DNA motifs. It is critical to optimize the DNA tokenization step towards the development of accurate, interpretable, and generalizable models.
@@ 2 Background
The k-mer and BPE-based tokenizers are commonly used in state-of-the-art DNA-LMs(Table 1). HyenaDNA (Nguyen et al., 2023) uses a 1-mer tokenizer and a decoder-only architecture with the Hyena operator to model long-range dependencies. DNABERT-1 (Ji et al., 2021) tokenized the genome with overlapping k-mers and trained separate models for each, substantially outperforming several downstream tasks. Nucleotide Transformer (Dalla-Torre et al., 2025) employs non-overlapping 6-mer tokenization and leverages large-scale pretraining across thousands of human individual genomes, and hundreds of species. Zhou et al. (2023) proposed DNABERT-2, which uses a byte-pair encoding (BPE) tokenizer inspired by NLP. MxDNA (Qiao et al., 2024) introduced a learnable tokenizer using a Mixture-of-Experts framework (Shazeer et al., 2017) and deformable convolutions (Dai et al., 2017).
@@
In addition, DNA is composed of highly repetitive, short, sparse, unevenly distributed, but conserved DNA sequence motifs across 600 million years of bilaterian evolution Nitta et al. (2015), largely represented by transcription factor (TF) binding motifs. The complexity of gene regulation often arises from the specific rearrangement and combination of these conserved motifs into context-specific regulatory elements Wong et al. (2020). Standard tokenization methods, which are agnostic to biological function, often arbitrarily split these meaningful motifs into smaller, non-functional tokens (see Figure 1b). As a result, it hampers DNA-LMs’ ability to learn biologically meaningful representations of genomic sequences and complicates downstream model interpretation.
@@ 3.1 Genomic sequences
Human reference genome: We use the most widely used human reference genome, version hg38 (GRCh38) (Consortium, 2013), which incorporates improved accuracy and coverage over previous releases.
Annotation of motif regions: We download the genome-wide JASPAR CORE TF motif predictions on the human reference genome (hg38) from the UCSC Genome Browser (Lee et al., 2020)(Raney et al., 2014). For BPE training, we extract all predicted motif sequences and merge any overlapping ones to create a non-redundant set. The resulting set of motif sequences covers approximately 59.84% (See Table C.1.1) of the human genome.
Annotation of cCREs: Candidate cis-regulatory elements (cCREs) are functional regulatory units in the genome, such as promoters and enhancers, that are typically hundreds of base pairs long. These regions often contain multiple TF motifs and are crucial for controlling when and where genes are expressed. For this study, we download human cCRE annotations from The Encyclopedia of DNA Elements(ENCODE), which provide genomic coordinates and regulatory classifications (Moore et al., 2020)(Luo et al., 2020). We extract the corresponding DNA sequences for BPE training, which constitute approximately 20.32% (See Table C.2.2) of the human reference genome.
@@ 3.2 Benchmark datasets
Genome Understanding Evaluation(GUE): Genome Understanding Evaluation (GUE) dataset was collected by (Zhou et al., 2023), consisting of 28 distinct datasets across 7 tasks and 4 species, with DNA inputs ranging from 70 to 1000 base pairs (bp). The metric we use is the Matthews Correlation Coefficient (MCC) (see Appendix).
Nucleotide Transformer Benchmarks: Nucleotide Transformer (NT) benchmarks were collected by (Dalla-Torre et al., 2025), it includes 18 datasets across 4 tasks only on humans. For this dataset, MCC is employed as the metric.
Dart-Eval: This dataset was introduced by (Patel et al., 2024) and contains five tasks. In our experiments, we focus on tasks 1–3. Accuracy (ACC) served as the evaluation metric. Task 1 involves distinguishing cCRE regions from background sequences. Task 2 requires identifying transcription factor (TF) binding motifs within background sequences. Task 3 entails classifying sequences specific to five different cell types against background sequences.
Genomic Benchmarks: This dataset was collected by (Grešová et al., 2023), it includes 9 tasks, across 9 species. We only use human related datasets. MCC is used as the metric.
SCREEN: We create this benchmark dataset by first downloading the cCREs on hg38 from the SCREEN interface (Moore et al., 2020), a platform for searching and visualizing the ENCODE Registry of candidate cis-Regulatory Elements (cCREs). This Registry contains 2,348,854 human cCREs, classified into eight categories, including promoter-like signatures (PLS), proximal and distal enhancer-like signatures (pELS, dELS), and CTCF- or TF–associated accessible elements (CA, CA-CTCF, CA-H3K4me3, CA-TF, TF). We generated a negative superset by taking the complement of all cCRE regions and dividing it into 300 base pair (bp) segments. For each cCRE category, we then randomly sample from this superset to obtain the same number of sequences as in the corresponding negative set. This procedure yielded eight datasets, each containing an equal number of positive and negative sequences.
@@ 4.1 Tokenizers
K-mer tokenizer: A k-mer is a substring of length k from a DNA or RNA sequence (Moeckel et al., 2024). K-mers can be generated in two ways: overlapping, by sliding a window one base at a time to capture all subsequences, or non-overlapping, by moving the window in steps of k to produce disjoint subsequences. In our experiments, we test overlapping 3-mer and 6-mer implemented by DNABERT-1 (Ji et al., 2021), and non-overlapping 6-mer implemented by Nucleotide Transformer (Dalla-Torre et al., 2025), respectively.
BPE tokenizer: BPE is a learnable subword tokenization algorithm that iteratively merges the most frequent pairs of characters or subwords in a training corpus to build a vocabulary of common subwords (Sennrich et al., 2015). The resulting merge rules, based on pair frequency, are recorded and applied to tokenize new sequences consistently, allowing the model to capture recurring patterns efficiently. The BPE tokenizer for DNA sequences was introduced by DNABERT-2. We use the BPE tokenizer pretrained by DNABERT-2 and also train our own BPEs on human reference genome, motif enriched genomic regions, and cis-regulatory element (cCRE) regions (see Section 3). Three vocabulary sizes were explored: 4096 (matching DNABERT-2), 2048, and 1024. Each BPE was initialized with the DNA alphabet (A, T, C, G, N) and five special tokens ([PAD], [UNK], [CLS], [SEP], [MASK]), with a minimum frequency threshold of 100.
Others: MxDNA (Qiao et al., 2024) is excluded from comparison due to a lack of publicly available pre-trained weights.
@@ 4.2 Models
Architecture: We adopt a BERT-based masked language model architecture (BertMLM) (Devlin et al., 2019) for our experiments. The model uses the Transformer encoder architecture (Vaswani et al., 2017) with 12 layers and 12 attention heads and a hidden size of 768, resulting in intermediate feed-forward layers of size 3072. The maximum input sequence length is 512 tokens, and the model uses learnable positional embeddings up to this length. Dropout Srivastava et al. (2014) is applied to both the attention probabilities and hidden layers with a rate of 0.1, and the GELU (Hendrycks & Gimpel, 2016) activation function is used throughout. The model vocabulary size is changed according to the tokenizer used, and a type vocabulary of size 2 is included to distinguish segment embeddings. All parameters are initialized with a standard deviation of 0.02.
@@
We compare the vocabulary similarities by calculating the pairwise Jaccard Index. We also generate Venn diagrams illustrating their word overlap (Hulsen, 2021). See Appendix for more details.
@@ 5 Benchmarking of State-of-the-art Tokenizers
We use the NT-benchmarks to investigate the zero-shot ability, where k-mer tokenizers consistently outperformed BPE tokenizers. Both overlapping and non-overlapping k-mer tokenizers struggled to cluster four distinct genomic elements, while BPE tokenizers performed better at separating enhancers and promoters from other DNA sequences (Figure 2c). Furthermore, we leverage the most comprehensive GUE benchmark and group downstream tasks by the length of DNA sequences. We find that k-mer tokenizers performed marginally better on shorter sequences (100 bp), whereas BPE tokenizers excel on longer ones (>>500 bp) (Figure 2d), which is consistent with previous work (Zhou et al., 2023). Taken together, we conclude that BPE-based tokenizers achieve more robust performance on both zero-shot and fine-tuning tasks and are more generalizable for predictions involving longer DNA sequences.
@@ 7.1 Algorithm
Motif processing: Motifs are short, recurring patterns in DNA, RNA, or protein sequences that are frequently associated with specific biological functions. In this work, we focus on TF motifs, which are short DNA sequences that are bound by transcription factor proteins to regulate gene expression. The TF motifs are generally represented in the form of position weight matrices (PWM)(Stormo, 2000), which indicate the probability of each nucleotide occurring at each position within the motif. We download motif PWMs from the JASPAR 2024 motif library (Sandelin et al., 2004)(Rauluseviciute et al., 2024), which is a widely used, open-access repository. We use the vertebrate library, which contains 879 non-redundant motifs. As most TF motifs range in length from 5 to 12 base pairs (bp), we exclude any motifs longer than 12 bp from our analysis. To incorporate TF motifs into our vocabulary, we binarize their PWMs and encode them into fixed sequences (Figure  5a). For each TF PWM, we apply a threshold probability of 0.5. Positions with lower probability are defined as wildcard positions. Subsequently, we discard wildcard positions at both ends of the motif. For the remaining positions, we encode them using the nucleotide with the highest probability.
@@ 7.2 Evaluation performance
Figure 6: Interpretation of DNAMotifTokenizer. (a) Workflow for the multiclass cell-type ATAC-seq classification task. (b) Distribution of average attribution scores for k-mer and motif tokens on the test set (10%). The motif tokens exhibit the strongest influence on model predictions. (c) Overlap between top 200 highest-attribution tokens from three cell types. (d) Highly attributed motif tokens matched with motif enrichment results from  Li et al. (2023).
Interpretability: We collected single-nucleus ATAC-seq (snATAC-seq) data generated from three diverse brain cell types: intratelencephalic neurons from cortical layer 2/3 (ITL23), VIP-positive GABAergic neurons (VIP), and Microglia (MGC) Li et al. (2023). We identified cell-type-specific ATAC-seq peak regions, which serve as prediction targets (Figure  6a). To control the vocabulary size and achieve fair comparison, we then fine-tuned our DNAMotifTokenizer models on this dataset. Then, we performed token attribution analysis using Integrated Gradients Sundararajan et al. (2017) on the test set (10%). For each token, we computed its attribution score toward the true class label and then averaged the scores; Tokens with attribution scores farther away from zero—either positive or negative—indicate stronger influence on the model’s prediction (Figure  6b). By ranking them, we found that in our DNAMotifTokenizer model, most of the highly influential tokens correspond to motif tokens (in red), whereas the k-mer tokens (in blue) generally had attribution scores close to zero. We selected the top 200 most contributive motif tokens for each cell type. The Venn plot indicates that DNAMotifTokenizer not only utilizes 143 motif tokens shared between three cell types, but also uses 25, 26, 34 cell-type-specific motif tokens for prediction in ITL23, VIP, MGC, respectively (Figure  6c). In addition, we compared these most contributive motif tokens with the enriched motifs for each cell type reported in the original paper Li et al. (2023). We showed top-matched motifs from three cell types as a heatmap (Figure  6d). In ITL23 cell type, motifs from the WT1, RORa, Egr, and KLF families are captured and enriched, whereas only WT1, and Egr2 are captured in the VIP cell type. In MGC cell type, SOX family motifs are captured and enriched. These results demonstrate the interpretability of motif tokens introduced by DNAMotifTokenizer in various cell types.